A country with a civilian population of 120,000 (all over age 16) has 100,000 employed and 10,000 unemployed persons, of which 5,000 are frictionally unemployed and another 3,000 structurally unemployed. Given the above data, complete the following questions, show your work: (20 points) a. What is the size of the labor force? b. What is the actual unemployment rate? c. What is the labor force participation rate? d. What is the employment/population ratio? e. What is the natural rate of unemployment?


1
Expert's answer
2021-06-07T13:09:42-0400

a)

Labor force = no. of employed + no. of unemployed people in the economy

Size of the labor force = 100,000 + 10,000

Size of the labor force = 110,000 people.


b)

 the actual unemployment rate

 = (No. of unemployed people ÷ Size of the       labor force) * 100

 = (10,000 ÷ 110,000) * 100

 = (0.0909) *100

 = 9.09%


c)

 = (size of the labor force ÷ Size of the population) * 100

  = (110,000 ÷ 120,000) * 100

  = (0.9166) * 100

  = 91.67%


d)

The employment/population ratio:

"\frac{Employment}{population}"


"=\frac{100,000}{120,000}=0.8333"


"=83.33\%"


e)

The natural rate of unemployment:

The rate of natural unemployment only includes structurally unemployed people.


"NRU = \frac{3000
}{108000}"


NRU = 2.78%